40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 14 00:42:40 2012, Jenne, Update, Locking, POX signal sometimes looks very funny POX_13Aug2012_LSCdisabled_pox_is_normalish.pngPOX_13Aug2012_LSCenabled_pox_is_squareish_funny.png
    Reply  Tue Aug 14 04:32:49 2012, rana, yoichi, Update, Locking, POX signal sometimes looks very funny ITMX_1028974969.bmpETMXF_1028975007.bmpETMXT_1028975105.bmpAS_1028975166.bmp
       Reply  Tue Aug 14 14:26:40 2012, Steve, Update, Cameras, cameras touched up cameras.png
          Reply  Fri Aug 17 08:33:46 2012, Steve, Update, Cameras, video cameras in the dark dark.jpg
             Reply  Fri Aug 17 09:34:27 2012, Koji, Update, Cameras, video cameras in the dark 
                Reply  Fri Aug 17 10:38:15 2012, Steve, Update, Cameras, video cameras in the dark 
             Reply  Fri Aug 17 14:45:51 2012, rana, Update, Cameras, video cameras in the dark 
                Reply  Mon Aug 20 09:49:01 2012, Steve, Update, Cameras, video cameras in the DARK onlyOSEMs.jpgOSEMsand_IR.jpgOSEMsandOplev633.jpg
Message ID: 7232     Entry time: Mon Aug 20 09:49:01 2012     In reply to: 7219
Author: Steve 
Type: Update 
Category: Cameras 
Subject: video cameras in the DARK 

Quote:

 The problem with the glow on the ETMY face is due to the red light being scattered off of the optical table from the HeNe laser for the OL. Why is the red light hitting the table?

One way to fix the problem for the camera image is to insert a long pass filter (if Steve can find one).

 Edmund Optics: NT62-874

 Edmund Optics: NT65-731

Edmund Optics: NT32-759 

 

 Atm1, condition: all oplev lasers are off or blocked, green shutters are closed at the ends, PSL out put shutter is closed, all outside LED illuminating are off, all room lights are off

                         Only the OSEMs are on. ETMY and ITMX are still look like illuminated.

Atm2, condition: open PSL shutter. ETMY at 11 o'clock  and ETMX 1 o'clock bright scattered spot of 1064 nm are visible

Atm3, condition: closed PSL shutter and restored all oplev He/Ne lasers, it is visible at ETMY

Next: I will disconnect power to OSEMs at ETMY

Attachment 1: onlyOSEMs.jpg  329 kB  | Hide | Hide all
onlyOSEMs.jpg
Attachment 2: OSEMsand_IR.jpg  324 kB  | Hide | Hide all
OSEMsand_IR.jpg
Attachment 3: OSEMsandOplev633.jpg  325 kB  | Hide | Hide all
OSEMsandOplev633.jpg
ELOG V3.1.3-